# LCSim: A Large-Scale Controllable Traffic Simulator

### v1

@@ Abstract
With the rapid development of urban transportation and the continuous advancement in autonomous vehicles, the demand for safely and efficiently testing autonomous driving and traffic optimization algorithms arises, which needs accurate modeling of large-scale urban traffic scenarios. Existing traffic simulation systems encounter two significant limitations. Firstly, they often rely on open-source datasets or manually crafted maps, constraining the scale of simulations. Secondly, vehicle models within these systems tend to be either oversimplified or lack controllability, compromising the authenticity and diversity of the simulations. In this paper, we propose LCSim, a large-scale controllable traffic simulator. LCSim provides map tools for constructing unified high-definition map (HD map) descriptions from open-source datasets including Waymo and Argoverse or publicly available data sources like OpenStreetMap to scale up the simulation scenarios. Also, we integrate diffusion-based traffic simulation into the simulator for realistic and controllable microscopic traffic flow modeling. By leveraging these features, LCSim provides realistic and diverse virtual traffic environments. Code and Demos are available at https://github.com/tsinghua-fib-lab/LCSim.
1 Introduction
@@
As global urbanization progresses, the complexity and diversity of urban transportation systems continue to increase. The driving styles of vehicles in different cities often have distinct characteristics [30], leading to high costs associated with testing and deploying related algorithms. For algorithms such as traffic optimization and autonomous driving, which require high levels of safety and reliability, thorough testing and evaluation before actual deployment are crucial [38]. This necessitates accurately modeling urban microscopic traffic flow through traffic simulation, enabling the safe and efficient assessment of relevant algorithms. Accurate modeling of urban traffic flow poses two main challenges for simulation systems: the need for realistic and controllable vehicle models to replicate the complex and diverse driving behaviors in reality; and the requirement for sufficiently realistic large-scale road network data to support traffic simulation.
@@
Existing simulation methods often have shortcomings in these two aspects. On one hand, vehicle modeling in simulation systems are typically categorized into three types: log-replay [13, 9, 23], rule-based (e.g. IDM [6]) simulation [2, 22, 39, 45, 13, 9, 23, 24, 46], and learning-based methods [1, 16, 10, 3, 15, 7, 50]. Log-replay lacks interaction, making it unable to create a closed-loop testing environment for vehicles. Rule-based simulations are often too simplistic and fail to replicate real vehicle behaviors accurately, while learning-based simulations usually lack controllability and struggle to model different driving styles accurately. CTG [50] has proposed a controllable traffic simulation method based on diffusion, but its scene is limited to the nuScenes [8] dataset, and the simulation environment has not been made open-source. On the other hand, many simulation systems rely on public datasets that only contain fragmented scenarios [9, 13, 23], limiting the scale of simulations. Metadrive [22] offers manual map creation tools, but this method is challenging to apply to large-scale urban scenarios. SUMO [2] provides map construction tools based on public data sources like OpenStreetMap (OSM) [^1] https://www.openstreetmap.org/, but its modeling of traffic flow is overly simplistic, extra efforts are needed to improve map formats and enhance vehicle modeling.
We propose LCSim, a Large-scale, Controllable traffic Simulator to address the abovementioned challenges. Our contributions are listed below: We provide an automated HD map generation and vehicle OD (origin-destination) flow process based on the MOSS toolchain [46]. This innovative approach enables researchers to reconstruct traffic scenarios from open-source datasets or create HD maps directly from public data sources such as OSM, employing latitude and longitude region delineation. By enhancing the scalability of simulations, our methodology offers new possibilities for analyzing complex traffic patterns. Our simulation framework, LCSim, incorporates a vehicle behavior model based on diffusion, allowing for realistic and controllable traffic simulations. By manipulating parameters such as vehicle target speed, maximum acceleration, and time headway, LCSim accurately captures diverse driving styles and characterizes traffic flows. We conducted a series of experiments to validate LCSim’s functionality. Firstly, we trained the diffusion model on the Waymo open dataset [26] and verified its performance on the validation set. Subsequently, we observed significant differences in vehicle behavior styles by comparing vehicle behavior data between Waymo and an internal dataset from SenseTime  [^2] https://www.sensetime.com/. We demonstrated that LCSim can simulate different vehicle styles by imposing control conditions on the diffusion process, accurately modeling the two distinct vehicle behavior patterns mentioned above. Moreover, we conducted training of single-agent reinforcement learning in different style simulation scenarios, showcasing the benefits of a multi-style simulation environment for training autonomous driving algorithms. Finally, we presented an application case of large-scale controllable traffic simulation based on LCSim.
2 Related Work
Traffic Simulators. The development of traffic simulators has a history of over a decade. Initially, researchers do simulations based on synthetic traffic scenarios and rule-based vehicle models [2, 12, 45, 24, 39]. However, these simulators’ rule-based models were often too simplistic, unable to model real and diverse vehicle behaviors accurately. With the continuous advancement of autonomous driving, an increasing number of open-source datasets containing real-world traffic scenarios have been released in recent years [26, 40, 8, 9, 14]. Consequently, many simulators based on these open-source datasets have emerged [20, 36, 22, 23, 13]. They utilize log-replay to rebuild realistic traffic scenarios and incorporate rule-based models to enable interactive simulations. Building upon this foundation, DriverGym [20] provides a learning-based vehicle model based on Simnet [3], while ScenarioNet [23] integrates various open-source datasets and offers reinforcement learning-based vehicle models. However, these traffic simulators based on open-source datasets often only simulate fragmented scenarios based on the provided data, and their simulation scale is limited to the scope of the dataset. Metadrive [22] presents a scenario-creation tool based on the combination of map elements. However, this approach faces challenges when it comes to constructing large-scale urban road networks. Furthermore, to our best knowledge, no open-source simulators currently provide controllable vehicle models to simulate multi-style vehicle behaviors in the real world.
Learning-based Traffic Simulation. Various learning-based vehicle simulation methods have emerged with the increasing availability of open-source traffic scenario datasets in recent years. Among them, imitation learning is often employed to learn expert actions from the dataset, thereby achieving realistic traffic simulation [41, 3, 5, 48, 4, 42]. However, this approach is often plagued by causal confusion [11] and distribution shift [29]. Reinforcement learning methods, on the other hand, address the distribution shift issue effectively by interacting with the simulation environment to learn driving behaviors [19, 16, 27, 37, 49]. However, the design of reward functions and the construction of the simulation environment are often complex. As generative models have advanced, many researchers have started to utilize the generation of vehicle behavior sequences for simulation purposes [33, 34, 47, 28, 35, 21, 50]. Among these approaches, CTG [50] utilizes a diffusion model to achieve controllable vehicle simulation. However, the code for CTG has not been made open-source. Following their idea, we trained our diffusion model based on the Waymo open dataset [26]. By generating vehicle behavior sequences, we were able to achieve controllable traffic simulation.
Table 1: Comparison of related traffic simulators. LCSim provides automated tools for map and origin-destination (OD) generation and controllable vehicle models, supporting multi-style traffic simulation within any given geographic region.
Map Tools
Long-term
@@ 2 Related Work
Controllability
SUMO [2]
nuPlan-devkit [9]
DriverGym [20]
Nocturne [36]
MetaDrive [22, 23]
Waymax [13]
Limsim [39]
TrafficSim [33]
SimNet [3]
CTG [50]
@@
3 System Design
Figure 1: The basic architecture of LCSim.
3.1 Diffusion Model
Scene Encoder. For accurately modeling the behavior of traffic participants, feature representations of scene information including map elements and historical states of traffic participants are required as conditions for the diffusion model. Following [51, 32], we utilize a spatial-temporal attention mechanism to model the scene features, taking in map polygons and historical states of compute scene embedding for each vehicle in the scenario.
Diffusion Process. We use the vehicle’s velocity and heading angle sequences as the generation target for the Diffusion model. Similar to QCNet [51], we employ an attention-based architecture for the diff decoder, which is utilized in the denoising process. The decoder takes the input data with noise and Fourier embeddings of the noise level as query values and performs cross-attention with the scene embeddings, resulting in denoised data as the output. The training process of the diffusion model follows Nvidia’s EDM architecture [18].
Guide Functions. Similar to CTG [50], we apply a loss function to the denoised results and backpropagate the gradients to guide the generation process of the diffusion model. In our experiments, the control objectives include the vehicle’s maximum acceleration, target velocity, time headway, and relative distance to the preceding car during car-following, and generating adversarial behavior by controlling nearby vehicles to approach the current vehicle. More details about how these loss functions are calculated can be found in the appendix.
3.2 Simulation Architecture
@@
LCSim performs discrete-time simulation based on a given time interval. Figure 1 illustrates the basic components of a simulation step. Each simulation step can be divided into two stages:
@@
  • Prepare Stage: During this stage, the simulator prepares the observation data for each vehicle, as depicted by the blue box in Figure 1. The observation data comprises three components: scene information observed by the vehicle, including road network topology and surrounding vehicles; vehicle embedding information computed by the scene encoder of the diffusion model, which encodes the scene information; and vehicles’ future action sequences either generated by the diffusion model conditioned on the scene encoding and guide function or given from the driving logs from open source dataset. Our diffusion model is trained on the Waymo dataset[26] and can generate vehicle action sequences for the future 8 seconds. During the simulation process, we employ a recurrent generation approach based on a specified step interval.
  • Update Stage: In this stage, each vehicle’s action is calculated by its policy based on the observation data, and these actions are used to update the vehicles’ states.
@@
We have implemented five different policies to support traffic simulation in various scenarios:
@@
  • RL-based Policy: A PPO [31] agent trained based on our simulator, its observation space contains the scene embedding and the action sequence. The action space consists of acceleration and steering values.
More details about action calculation and parameters of each policy can be found in the appendix.
3.3 Map & OD Generator
Figure 2: The process of map generation.
We have defined a unified map and vehicle Origin-Destination (OD) format based on Protobuf. Additionally, we have developed format conversion tools designed for the Waymo and Argoverse datasets. Our map and OD generator were implemented based on the MOSS toolchain [46]. Figure 2 illustrates the process of generating an HD map based on OSM. We crawl OSM data in GeoJSON format within a specified latitude and longitude range. This data includes road network information. We then complete the topological connections between road segments and add lane markings and drivable areas based on traffic rules. The final output is an HD map following the unified format we defined.
For OD generation, We provide a random vehicle OD generation tool similar to SUMO [2]. Additionally, we have integrated the OD generation algorithm based on deep generative models from the MOSS [46] toolchain. This algorithm leverages deep learning techniques to generate OD pairs that closely resemble real-world traffic patterns.
4 Experiments
We conducted a series of experiments to validate LCSim’s functionality. First We show how the diffusion model within LCSim is constructed and its ability to conduct realistic and controllable traffic simulations, then we conducted reinforcement learning training within different styles of scenarios generated by LCSim.
4.1 Diffusive Simulation Construction
Datasets. We trained our diffusion model on the Waymo Open Motion Dataset [26], which contains 500+ hours of driving logs collected from seven different cities in the United States. The dataset is further divided into scene segments of 20s and 9s. In this experiment, we utilized the 9s segments for training, using the initial 1s as the historical context, and let the model generate future 8s sequences of vehicle behaviors. Subsequently, we utilized the generated data for traffic simulation. The diffusion model was trained on the training set and evaluated for its simulated performance on the validation set.
Metrics & Settings. We propose two evaluation metrics to validate the Diffusive Simulation’s authenticity. The generated vehicle behavior sequences must first adhere to the fundamental traffic rules. Within the 8s simulated scenarios, we calculate the probabilities of collision occurrence and vehicles driving off the road. These probabilities are then compared to the corresponding ones derived from the real log data in the Waymo dataset. In this calculation process, we filter out vehicles parking outside the boundaries of the road in the scene. Secondly, to assess the fidelity of the simulation, we compare the distribution of vehicle behaviors in both the real dataset and the simulated data. We compute the Kullback-Leibler (KL) divergence between these distributions. The vehicle behaviors considered in this analysis include acceleration, velocity, and time headway (THW). In the calculation process, we filter out vehicles that remain static throughout the entire scene segment. The time headway is only taken into account when the current vehicle is detected to be engaged in a car-following behavior. In both the simulation based on the Waymo driving log and the diffusion model, we employed the BicycleExpertPolicy to ensure that vehicle behavior adhered to kinematic constraints, thus mitigating the impact of noise in trajectory data. To handle the impact of randomness in the generation process, we conducted repeated experiments and reported the mean and error bars of the results. The calculation of probability distributions utilized all the data generated from the repeated experiments. More details of the experiment can be found in the appendix.
Table 2: Evaluation results of diffusive traffic simulation.
Collision (%) Off-Road (%) KL-Divergence
ACC VEL THW
Waymo 2.664 ($\pm$ 0.000) 1.521 ($\pm$ 0.000) -
Diffusive Sim 2.916 ($\pm$ 0.160) 1.602 ($\pm$ 0.112) 0.0197 0.0136 0.0465
Figure 3: The comparison of vehicle behaviors in the Waymo dataset and diffusive simulation.
Results. Table 2 presents the quantitative results, which shows that the simulated results in terms of collision and off-road probabilities closely align with the statistics derived from the Waymo dataset. This indicates that our diffusion model, based on its ability to learn from real-world vehicle behaviors, successfully captures the fundamental traffic rules that govern vehicle movements. Regarding the distribution of vehicle behaviors, as depicted in Figure 3, the consistency between the simulated vehicle behavior distributions generated by our model and the statistical results from real driving logs can be observed. This consistency highlights the model’s ability to learn the characteristics of vehicles in the dataset. Taken together, these findings indicate that our diffusion model is capable of conducting realistic traffic simulations based on real-world scenarios.
4.2 Aligning Vehicle Behavior Characteristics
SenseTime Driving Dataset. SenseTime driving dataset comprises about 400 hours of vehicle driving logs collected from vehicles based on SenseAuto [^3] https://www.senseauto.com/ in the Beijing Yizhuang area. The data is presented in a format similar to vehicle trajectories in the Waymo dataset. We conducted statistical analysis on the dataset, focusing on metrics such as acceleration, relative distance, and time headway during the car following process. This analysis allowed us to derive the driving behavior characteristics of vehicles in the Yizhuang area.
In Figure 4, we compared the behavioral characteristics of vehicles in the SenseTime dataset with those in the Waymo dataset. The comparison included metrics such as vehicle acceleration, relative distance, and time headway during car following. Since vehicle driving speeds are often related to the specific driving environment (e.g., road congestion, lane speed limits) rather than the behavioral characteristics of vehicles, we did not include speed in the comparison. It can be observed that there are significant differences in the behavioral characteristics of vehicles between these two datasets collected from different regions. Vehicles in the Yizhuang area exhibit a more "gentle" driving style, showing a preference for using smaller accelerations during acceleration and braking processes. Additionally, they maintain larger relative distances and headway times during the following process compared to vehicles in the Waymo dataset.
Figure 4: The differences in vehicle behavior between Waymo and SenseTime datasets.
As our diffusion model is trained on the Waymo dataset, without imposing any control conditions on the generation process, the vehicle behavior characteristics in the simulation remain consistent with the Waymo dataset, as shown in Figure 3. By applying constraints such as max acceleration, relative distance, and time headway during car-following in the generation process, we can align the vehicle behavior characteristics produced by the diffusive simulation with those collected in the SenseTime driving dataset. The comparison of the two distributions can be seen in Figure 5. This demonstrates that our simulator can model vehicles with diverse driving styles, thereby providing different styles of traffic simulation environments. Details about the guide function we use here can be found in the appendix.
Figure 5: The comparison of vehicle behaviors in the SenseTime dataset and guided diffusive simulation.
4.3 Multi-Style RL Training
We constructed a single-agent reinforcement learning environment based on the Waymo dataset with our guided diffusive simulation to further investigate the impact of traffic environments with different driving styles on driving policy learning. Below are the training settings and results:
Settings. To validate the simulation model’s capability in unseen scenarios, we constructed a reinforcement learning environment based on the validation set of the Waymo dataset. We selected 4,400 scenarios from the validation set and further divided them into a training set containing 4,000 scenarios and a test set containing 400 scenarios. We trained a PPO [31] agent on the training set and evaluated its performance on the test set. The agent’s observation space consisted of scene embedding generated by the diffusion model’s scene encoder and the vehicle’s reference route. The route was either based on real trajectories from the driving logs or accumulated action sequences generated by the diffusion model in different training environments. On the test set, we tested four metrics of the agent: collision rate, off-road rate, average route progress rate (We project the current position of vehicles onto the Frenet coordinates of the route and calculate the ratio of the completed length to the total length.), and scene success rate (i.e., successfully reaching the destination without collisions or driving off-road). We also provided the average reward value per episode. The vehicle behaviors of the test set are based on Waymo driving logs. Detailed RL training settings can be found in the appendix.
Styles of Training Environments. We created four distinct driving environments on the training set: In the first environment, vehicles base their actions on real trajectories from the Waymo driving logs. The second environment utilizes the diffusion model without guide functions, which maintains consistency with the Waymo dataset in terms of vehicle behavior styles. With the diffusion model’s nature, it generates diverse vehicle trajectories under the same initial conditions, exposing the agent to a broader range of traffic scenarios during training. The third environment follows the driving style observed in the SenseTime driving dataset, emphasizing a more "gentle" driving behavior compared to the Waymo-based environment. Furthermore, an adversarial driving environment is implemented by guiding nearby vehicles closer to the agent, creating a training scenario with a higher potential for collisions. Details about the configuration of guiding functions for each environment are available in the appendix.
Table 3: Evaluation results of RL agents.
Waymo 15.71 ($\pm$ 0.56) 4.24 ($\pm$ 0.20) 84.98 ($\pm$ 0.53) 51.31 ($\pm$ 0.56) 7.41 ($\pm$ 0.10)
Diff 12.06 ($\pm$ 0.28) 3.01 ($\pm$ 0.12) 84.76 ($\pm$ 0.35) 52.75 ($\pm$ 0.50) 8.59 ($\pm$ 0.14)
Results. As shown in Table 3, the results show that the agent trained in the diffusive simulation environment with the same driving style as Waymo outperformed the agent trained based on Waymo driving logs across various metrics. This suggests that the diffusion model can generate diverse traffic scenarios based on initial conditions, contributing to the learning of driving policy. The performance of the agent trained in the environment using the driving style of the SenseTime dataset was not good in relatively more aggressive environments. The difference between the learned behavior style of the agent and the testing environment led to an increase in collision rate, and the agent’s tendency towards a more "gentle" behavior style compared to the given route also led to a low route progress rate. On the other hand, training the agent in a more adversarial environment resulted in a lower collision rate. However, the agent’s proactive avoidance of nearby vehicles led to an increase in the off-road rate.
4.4 City-Scale Traffic Flow Construction
Figure 6: City-scale traffic flow simulations in Jinan and Shenzhen.
We showcased the scalability of LCSim with city-scale simulations on real-world traffic flow in two metropolises.
Datasets. We used two vehicle trajectory datasets[44, 43, 25] each with one-day-long city-scale trajectories of the entire fleet recovered from daily urban traffic camera videos in Jinan and Shenzhen city. Both of the dataset involves over one million trajectories and one thousand square kilometers of urban area.
Settings. Compared with GPS-based trajectories, the trajectories recovered from traffic cameras are temporally sparser, where only the arrival time at road intersection is specified [44], which is thus taken as the travel schedule of each agent in LCSim with a series of trips between intersections with corresponding departure time and arrival time. We showed that by simulating the driving behaviors of vehicles given their schedules, using the arrival time at a specific position as guidance, LCSim can effectively replicate real-world city-scale traffic conditions.
Results. Figure6 shows, in Jinan and Shenzhen, the spatial distribution of simulated road flow and speed, as well as the probability distribution of the arrival time error which is the deviation of the simulated arrival time of each trip compared with that of the ground truth trajectory data. As can be seen, the arrival time error is mainly distributed around zero with over 90% of trips having arrival time errors less than 20 seconds. LCSim also produced reasonable traffic conditions with coherence between the road network structure, flow, and speed.
5 Conclusion
We proposed LCSim, a large-scale, controllable diffusion-based traffic simulator. With an automated HD map generation and vehicle OD flow process, LCSim is capable of conducting large-scale traffic simulations on any selected region. With the integration of the diffusion model and guide functions, LCSim can simulate different styles of vehicle behaviors.
Limitations. LCSim has two main limitations. Firstly, the simulator is implemented in Python using a single-threaded CPU, which limits its performance potential. Although parallel simulation using multiple processes is currently employed as a solution, it does not fundamentally address the issue. One potential approach to overcome this limitation is to develop a multi-threaded version of the simulator using C++ and deploy the Diffusion model in C++, which is a potential direction for future work. Secondly, the simulator currently provides visualization only from a top-down perspective and lacks the rendering of realistic perceptual data. Integrating image generation based on Diffusion into the simulation is one of the planned future developments to address this limitation.
Appendix
In the appendix, we provide more details about the experiments discussed in the main text. Section  A introduces the implementation details of the diffusion model and the specific content and form of the guide function. Section B details the implementation of the system and showcases the visualization of scenarios in the simulator. Section C covers the relevant content of the SenseTime driving dataset, while Section D delves into the detailed experimental configurations for multi-style reinforcement learning experiments. Code and Demos are available at https://github.com/tsinghua-fib-lab/LCSim.
Appendix A Diffusion Model
The process of the diffusion model generating vehicle action sequences is shown in Figure 7. With the road network topology and vehicle historical states as input, the model generates future action sequences for the vehicle through a denoising diffusion process.
Due to the relevant regulations of the Waymo Open Motion Dataset (WOMD) [26], we cannot provide the parameters of the model trained on it. In this section, we introduce the implementation details of the diffusion model and the hyperparameters used for training and inference in detail to ensure that the relevant experimental results can be easily reproduced.
Figure 7: The process of generating vehicle action sequences by diffusion model.
A.1 Problem Formulation

### v2

@@ Abstract
With the rapid growth of urban transportation and the continuous progress in autonomous driving, a demand for robust benchmarking autonomous driving algorithms has emerged, calling for accurate modeling of large-scale urban traffic scenarios with diverse vehicle driving styles. Traditional traffic simulators, such as SUMO, often depend on hand-crafted scenarios and rule-based models, where vehicle actions are limited to speed adjustment and lane changes, making it difficult for them to create realistic traffic environments. In recent years, real-world traffic scenario datasets have been developed alongside advancements in autonomous driving, facilitating the rise of data-driven simulators and learning-based simulation methods. However, current data-driven simulators are often restricted to replicating the traffic scenarios and driving styles within the datasets they rely on, limiting their ability to model multi-style driving behaviors observed in the real world. We propose LCSim, a large-scale controllable traffic simulator. First, we define a unified data format for traffic scenarios and provide tools to construct them from multiple data sources, enabling large-scale traffic simulation. Furthermore, we integrate a diffusion-based vehicle motion planner into LCSim to facilitate realistic and diverse vehicle modeling. Under specific guidance, this allows for the creation of traffic scenarios that reflect various driving styles. Leveraging these features, LCSim can provide large-scale, realistic, and controllable virtual traffic environments. Codes and demos are available at https://tsinghua-fib-lab.github.io/LCSim.
Index Terms:
Traffic Simulation, autonomous driving, diffusion model.
I Introduction
@@
As global urbanization progresses, the complexity and diversity of urban transportation systems continue to increase. The driving styles of vehicles in various cities often have distinct characteristics [1], leading to high costs of benchmarking autonomous driving algorithms, which require high levels of safety and reliability and thorough testing and evaluation before actual deployment [2, 3]. This necessitates accurately modeling urban microscopic traffic scenarios through traffic simulation, enabling the robust assessment of relevant algorithms. Accurate modeling of urban traffic scenarios poses two main challenges for simulation systems: the need for realistic and controllable vehicle models to replicate the complex and diverse driving behaviors in reality, and the requirement for large-scale traffic scenario data to support traffic simulation.
@@
Existing simulation methods often have shortcomings in these two aspects. For one, vehicle behavior modeling in simulation systems is typically categorized into three types: rule-based (e.g. IDM [4]) simulation [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], log-replay of real-world dataset [9, 10, 11], and learning-based methods [15, 16, 17, 18, 19, 20, 21, 22]. Rule-based simulations are often too simplistic and fail to replicate real vehicle behaviors accurately. Log replay and learning-based methods excel in replicating real vehicle behavior, but they usually lack controllability and struggle to model different driving styles faithfully. CTG [21] has proposed a controllable traffic simulation method based on a diffusion model. But, its scenario is limited to the nuScenes [23] dataset, and this method has not been integrated into a simulation system for algorithm benchmarking. On the other hand, most data-driven simulators rely on public datasets that only contain fragmented scenarios [10, 9, 11], limiting the scale of the simulation. Metadrive [6] offers manual map creation tools. ScenarioNet [11] does a great job of collecting large-scale real-world traffic scenarios from various driving datasets. However, the driving styles in the simulation environments they provide are still constrained by the given dataset. Therefore, extra efforts are needed to improve traffic scenario construction and enhance vehicle modeling.
We propose LCSim, a Large-scale, Controllable traffic Simulator to address the abovementioned challenges. Our contributions are listed below:
  • We define a unified data format for traffic scenarios and provide tools to construct them from multiple data sources including real-world driving datasets like the Waymo open motion dataset (WOMD) [24] and Argoverse [25] dataset, and hand-crafted scenarios built from public data sources such as OpenStreetMap (OSM) [^1] https://www.openstreetmap.org/ [5, 13].
  • We design and implement a simulation system that integrates a diffusion-based vehicle motion planner to achieve realistic, diverse, and controllable traffic simulation in the constructed traffic scenarios. A Gym-like environment interface is provided to support reinforcement learning algorithm training and benchmarking.
  • A series of experiments are conducted to validate LCSim’s functionality. Firstly, we demonstrate the ability of the diffusion-based motion planner on WOMD. Next, reinforcement learning agents are trained in environments with different driving styles built by LCSim, showcasing the impact of various driving styles on algorithm benchmarking. Lastly, through the accurate replication of city-level traffic scenarios, we highlight LCSim’s capability to construct large-scale traffic simulations.
II Related Work
TABLE I: Comparison of related traffic simulators. LCSim provides automated tools for traffic scenario construction and diffusion-based controllable vehicle motion planning.
Scenario
Construction
RL
Environment
@@ II Related Work
Controllable
Sensor Sim
SUMO
nuPlan-devkit
DriverGym
MetaDrive
Waymax
TrafficSim
SimNet
CTG
@@
II-A Traffic Simulators
The development of traffic simulators has a history of over a decade. Initially, researchers conducted simulations based on hand-crafted traffic scenarios and rule-based vehicle models [5, 26, 8, 12, 7]. However, these rule-based models are often simplistic and unable to accurately model real and diverse vehicle behaviors. With the continuous advancement of autonomous driving, an increasing number of open-source datasets containing real-world traffic scenarios have been released in recent years [24, 25, 23, 10, 27]. Consequently, many data-driven simulators based on these datasets have emerged [28, 29, 11, 9]. They utilize log replay to rebuild realistic traffic scenarios and incorporate rule-based models to enable close-looped simulations. Building upon this foundation, DriverGym [28] provides a learning-based vehicle model based on SimNet [18], while ScenarioNet [11] integrates various open-source datasets and offers reinforcement learning-based vehicle agent. However, these data-driven simulators often only simulate fragmented scenarios based on the provided data, and their simulation scale is limited only to the scope of the dataset. Metadrive [6] presents a scenario-creation tool based on the combination of map elements. However, this approach faces challenges when it comes to constructing large-scale urban road networks. Furthermore, to the best of our knowledge, LCSim is the first open-source traffic simulator to provide controllable learning-based vehicle models to simulate multi-style driving behaviors in the real world.
II-B Learning-based Traffic Simulation
Various learning-based vehicle simulation methods have emerged with the increasing availability of open-source traffic scenario datasets in recent years. Among them, imitation learning is often employed to learn expert actions from the dataset, thereby achieving realistic traffic simulation [30, 18, 31, 32, 33, 34, 35, 36]. However, this approach is often plagued by causal confusion [37] and distribution shift [38]. Reinforcement learning methods, on the other hand, address the distribution shift issue effectively by interacting with the simulation environment to learn driving behaviors [39, 16, 40, 41, 42]. However, the design of reward functions and the construction of the simulation environment are often complex. As generative models have advanced, many researchers have started to utilize the generation of vehicle motion plans for simulation purposes [43, 44, 45, 46, 47, 48, 21]. Among these approaches, CTG [21] utilizes a diffusion model to achieve controllable vehicle simulation. However, this method has not been used in traffic simulators for further algorithm training and testing. Following their idea, we trained our diffusion-based vehicle motion planner based on WOMD. Controllable traffic simulation can be achieved by generating vehicle motion plans with guide functions.
III System Design
III-A Scenario Data Construction
To achieve large-scale traffic simulation, we define a unified data format based on Protobuf for traffic scenario data from multiple data sources including real-world driving datasets and hand-crafted scenarios provided by MOSS [13] toolchain.
Figure 1: The unified format of traffic scenario data.
III-A1 Unified Scenario Data Format
Figure 1 shows an example of our unified scenario data format. Each scenario data consists of the following two parts:
  • Map: The map data consists of three components: lanes, roads, and junctions. The lanes contain the primary map information, with each lane element storing the lane’s ID, type, polygon information of the centerline, and the connections between the current lane and others (such as predecessors, successors, and neighbors). Roads and junctions serve as containers for lane elements, both containing boundary information for the drivable areas on the map. Each lane belongs to a unique road or junction. Additionally, junctions store information related to traffic lights, including the IDs of the lanes they control and the traffic signal phases.
  • Agents: The Agents data contains all the information about agents that need simulating. Each agent element includes the agent’s ID, attributes, and routes. The attributes consist of basic properties such as the agent’s type and shape, while the routes contain the full travel schedules for the agent. An agent’s schedule data includes the departure time and reference route, with the reference route detailing the agent’s state (position, speed, heading, etc.) at each step after the departure time. In log replay-based simulations, this information is provided by scenario data, whereas in diffusion-based simulations, it is generated by the diffusion model and continuously updated throughout the simulation.
Figure 2: Traffic scenarios from different data sources.
III-A2 Scenario Construction
Figure 2 illustrates traffic scenarios constructed from different data sources. For the WOMD and Argoverse datasets, all map information is placed within a single junction. We align the basic attributes of the map features based on the map element types provided in WOMD. For the scenarios obtained from the MOSS toolchain, we complete the map elements, including drivable boundaries and road lines as the original data only contains centerlines. In the MOSS scenario, agent routes are provided as origin-destination points, and we implement an A-star-based router to complete them into full reference routes. MOSS allows the construction of arbitrarily large scenarios using latitude and longitude ranges. We further divide the map into roads and junctions, and during the simulation, the map elements and agents in each road or junction are organized into a data instance. These instances are processed in batches by the diffusion model for vehicle motion planning, allowing users to select areas of interest by ID and enable diffusion-based simulation within those areas.
III-B Simulation Architecture
Figure 3: The simulation architecture of LCSim.
@@ III-B Simulation Architecture
With scenario data constructed from multiple sources, LCSim performs discrete-time simulation based on a given time interval. Figure 3 illustrates the basic components of a simulation step. Each simulation step can be divided into two stages:
III-B1 Prepare Stage
@@
During this stage, the simulator prepares the observation data for each vehicle, as depicted in the blue box in Figure 3. The observation data comprises three components: scene information observed by the vehicle, including road network topology and surrounding vehicles, scene embeddings computed by the scene encoder, which encodes the scene information, and vehicles’ motion plans either generated by the diffusion decoder or given by the logs from the scenario data.
III-B2 Update Stage
In this stage, each vehicle’s action is calculated by its policy based on the observation data, and these actions are used to update the vehicles’ states. We implement various control policies in the simulator to handle different simulation scenarios. The ExpertPolicy controls vehicles to strictly follow the given motion plans. At the same time, the BicycleExpertPolicy enhances this by adding kinematic control based on the bicycle model to achieve more realistic simulation effects. Furthermore, we implement the IDMPolicy to enable closed-loop traffic simulation, where vehicles adjust their accelerations based on the objects ahead while following the given motion plans. Additionally, any vehicle in the scenario can be controlled by external input actions, allowing the simulator to serve as a training or testing environment for specific vehicle control algorithms.
@@ III-B2 Update Stage
For details, we implement the following five different policies to support traffic simulation in various scenarios:
@@
  • RL-based Policy: A PPO [49] agent trained based on our simulator.
III-C Diffusion-based Motion Planner
We design and implement a vehicle motion planning module based on a diffusion model to achieve controllable traffic simulation. During the simulation process, this module takes the scene information of the current step and a guide function as inputs to generate realistic and controllable motion plans for vehicles in the scenario. Algorithm 1 summarizes the guided generation process of the model.